OpenStudy (anonymous):

write the equation of the line passing through the points (-2,5) and (4,9) show steps thanks guys need the help bad

OpenStudy (anonymous):

gradient: 2/3 y-y1=m(x-x1) y-9=2/3(x-4) y-9=2x/3 -8/3 y= 2x/3 + 6.33

OpenStudy (anonymous):

Non-decimal form; \[y = (2 x)/3+19/3\]

OpenStudy (anonymous):

thanks for the help guys!! probably will need a lot more was dumb an waited for the day before the final to finally start reviewing and stuff haha

OpenStudy (anonymous):

how did you get the gradient 2/3

OpenStudy (anonymous):

y2-y1/x2-x1 9-5/4+2

OpenStudy (anonymous):

thanks alot!

OpenStudy (anonymous):

9-5 would be 4 though not 3??

OpenStudy (anonymous):

yeah, so 4/6 reduced is 2/3

OpenStudy (anonymous):

did not notice that it was negative haha thanks again

OpenStudy (anonymous):

sorry that im asking this again but y1, would be 5 right??

OpenStudy (anonymous):

and if im wrong can you explain why please and thank you